How do I find an equation of the line containing (-1,-3) and (-4,-9)? Writing it as a linear function in slope-intercept form.
Answer by mananth(16946)

x1 y1 x2 y2
-1 -3 -4 -9

slope m = (y2-y1)/(x2-x1)
( -9 - -3 )/( -4 - -1 )
( -6 / -3 )
m= 2
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( -1 , -3 ) in
Y = m x + b
-3.00 = 2 * -1 + b
-3.00 = -2 + b
b= -3 + 2
b= -1
So the equation will be
Y = 2 x -1
m.ananth@hotmail.ca